OpenStudy (lukecrayonz):

Find the area of the triangle: B=72.5 degreees, a=105, c=64

OpenStudy (amistre64):



OpenStudy (amistre64):

105 sin(72.5) = height and 64 is your base

OpenStudy (amistre64):

areaTri = bh/2

OpenStudy (amistre64):

64 sin(72.5) is just as viable a height, and 105 a base so it really doesnt matter how you squish them up; just divide the results in half

OpenStudy (lukecrayonz):

So b=100~?

OpenStudy (amistre64):

it says to find the area of the trianlge; if you wanna find b thats another story tho

OpenStudy (amistre64):

i imagine the length of B is according to the law of cosines b^2 = a^2+c^2-2ac cos(72.5)

OpenStudy (lukecrayonz):

I'm just trying to find b so I can do herons formula.